Draw a graph of the line passing through the points(-6,-7) and(-3,-5).Does the
point (3,-1)also lie on this line?

Answer by nerdybill(7384)

Draw a graph of the line passing through the points(-6,-7) and(-3,-5).Does the
point (3,-1)also lie on this line?
.
Algebraically:
slope m = (-5+7)/(-3+6) = 2/3
.
y - y1 = m(x - x1)
y - (-5) = (2/3)(x - (-3))
y + 5 = (2/3)x + 2
y = (2/3)x - 3 (this is the equation of the line)
.
Testing the point:(3,-1)
y = (2/3)x - 3
-1 = (2/3)3 - 3
-1 = 2 - 3
-1 = -1
.
So, YES, the point does lie on the line.
.